# Rydberg Atomic Receiver: Next Frontier of Wireless Communications

### v2

@@ VI Future Research Directions
  • More accurate transmission model. The current signal transmission model of RARE mainly considers the interaction between Rydberg atoms and EM waves. In practice, the Rydberg-Rydberg interaction can significantly influence the dynamics of quantum states, which should be thoroughly studied to make the signal transmission model of RARE more accurate.
  • Joint transmitter and RARE optimization. As RARE changes the signal transmission model from linear to nonlinear, traditional signal processing methods become suboptimal. It is essential to develop new transmission techniques tailored for both the RARE and transmitter, including modulation, precoding, channel estimation, synchronization, interference mitigation, and so on.
  • Physical layer security. RARE also holds promise in enhancing physical layer security. Taking advantage of the abundant energy levels, a RARE can flexibly switch the frequency band it is listening to. This capability enables the implementation of high-speed and ultra-broadband frequency hopping communication, which can effectively protect the privacy of wireless data.
  • Satellite communication and low-altitude economy. Another promising application of RARE is to assist satellite communication and low-altitude economy. The high sensitivity of RARE not only enables the expansion of satellite signal coverage but also demonstrates the potential to enhance the precision of sensing and navigation for unmanned aerial vehicles.
@@ VII Conclusions
This article provides an overview of RARE-aided wireless communications. Representing a paradigm shift in receiver architecture, RAREs detect EM signals via quantum phenomena, contributing to improved sensitivity and broadened detecable frequency range. RAREs can seamlessly integrate with both classic techniques, such as FDM and MIMO, as well as quantum techniques, like the many-body effect, to enhance communication performance. The potential of RARE in multiband S&C systems is revealed through experimental results. Finally, we have highlighted several promising future research directions on RARE.

### v3

@@ VI Future Research Directions
  • More accurate transmission model. The current signal transmission model of RARE primarily considers the interaction between Rydberg atoms and EM waves. In practice, Rydberg-Rydberg interactions can significantly influence the dynamics of quantum states. A comprehensive study of these interactions is essential to enhance the accuracy of the RARE signal transmission model.
  • Joint transmitter and RARE optimization. As RARE modifies the signal transmission model from linear to nonlinear, traditional signal processing methods become suboptimal. It is essential to develop new transmission strategies tailored for both the RARE and transmitter, including modulation, precoding, channel estimation, synchronization, interference mitigation, and so on.
  • Holographic atomic-MIMO communications. Conventionally, the performance of holographic MIMO—where massive antennas are deployed within a compact area—is limited by the mutual coupling between antenna elements. Fortunately, the holographic atomic-MIMO technique can overcome this limitation, as Rydberg atoms in different laser streams minimally influence each other.
  • Data-driven RARE. To date, the quantum dynamics of Rydberg atoms under complicated EM environments, such as twisted radiowaves and broadband signals, remains largely unexplored due to the intricate Lindblad master equations. Data-drive methods, including physics-informed neural networks and foundation models, offer promising solutions for understanding and optimizing these RARE systems [10].
  • Physical layer security. RARE holds significant potential for enhancing physical layer security. Leveraging abundant energy levels, RARE systems can flexibly switch their listening frequency bands. This capability enables high-speed and ultra-broadband frequency-hopping communication, effectively protecting wireless data privacy.
  • Satellite communication and low-altitude economy. Another promising application of RARE is in supporting satellite communication and the low-altitude economy. The high sensitivity of RARE not only enables the expansion of satellite signal coverage but also demonstrates the potential to enhance the precision of sensing and navigation for unmanned aerial vehicles.
@@ VII Conclusions
This article has provided an overview of RARE-aided wireless communications. Representing a paradigm shift in receiver architecture, RAREs detect EM signals via quantum phenomena, contributing to improved sensitivity and broadened detectable frequency range. RAREs can seamlessly integrate with both classic techniques, such as FDM and MIMO, as well as quantum techniques, like the many-body effect, to enhance communication performance. The potential of RARE in MSAC systems was revealed through experimental results. Finally, we have highlighted several promising future research directions on RARE.
Acknowledgment
This work was supported in part by the National Natural Science Foundation of China (No. 624B2123), in part by the Research Grants Council of the Hong Kong Special Administrative Region, China under a fellowship award (HKU RFS2122-7S04), NSFC/RGC CRS (CRS_HKU702/24), the Areas of Excellence scheme grant (AoE/E-601/22-R), Collaborative Research Fund (C1009-22G), and the Grant 17212423, and in part by the Shenzhen-Hong Kong-Macau Technology Research Programme (Type C) (SGDX20230821091559018).